Matthew is 3 times as old as Jenny. In 7 years, he will be twice as old as she will be then. How old is each now?
Found 2 solutions by stanbon, robertb:
Answer by stanbon(75887)

Matthew is 3 times as old as Jenny. In 7 years, he will be twice as old as she will be then. How old is each now?
-----------
Equations:
m = 3j
---
m+7 = 2(j+7)
-----
Substitute for "m" and solve for "j":
3j+7 = 2j+14
j = 7 (Jenny's age now)
---
Since m = 3j, m = 21 (Matthew's age now)
=======================
Cheers,
Stan H.
===========

Answer by robertb(5830)

m = 3j,
m+7 = 2(j+7). Substituting the first equation into the 2nd, we get
3j+7 = 2j+14, and so
j = 7, and m = 21.